# ChainMPQ: Interleaved Text-Image Reasoning Chains for Mitigating Relation Hallucinations

### v1

@@ Abstract
While Large Vision-Language Models (LVLMs) achieve strong performance in multimodal tasks, hallucinations continue to hinder their reliability. Among the three categories of hallucinations, which include object, attribute, and relation, relation hallucinations account for the largest proportion but have received the least attention. To address this issue, we propose ChainMPQ (Multi-Perspective Questions guided Interleaved Chain of Image and Text), a training-free method that improves relational inference in LVLMs by utilizing accumulated textual and visual memories. ChainMPQ first extracts subject and object keywords from the question to enhance the corresponding image regions. It then constructs multi-perspective questions that focus on the three core components of a relationship: the subject, the object, and the relation that links them. These questions are sequentially input to the model, with textual and visual memories from earlier steps providing supporting context for subsequent ones, thereby forming an interleaved chain of images and text that guides progressive relational reasoning. Experiments on multiple LVLMs and benchmarks show that ChainMPQ substantially reduces relation hallucinations, while ablation studies further validate the effectiveness of its three core modules.
@@ 1 Introduction
Although prior work has substantially reduced object and attribute hallucinations through preference optimization Wu et al. (2025a); Yang et al. (2025), contrastive decoding Wu et al. (2025b); Zhang et al. (2025), and intermediate-layer modification Jiang et al. (2024; 2025); Wang et al. (2024), relation hallucination, which accounts for nearly 40% of all hallucinations Zheng et al. (2024), remains a salient challenge. Existing efforts to mitigate relation hallucinations have made some progress: dataset-driven approaches Xie et al. (2025) focus on constructing high-quality training and fine-tuning data, prompt-engineering methods Wu et al. (2025a) employ constrained perception prompts and Detect-then-Calibrate Zheng et al. (2024) selectively adjust the final logits with intermediate-layer.
Despite their promising results, existing methods targted at relation hallucinations treat relational reasoning as single-step inference, expecting models to simultaneously identify entities and determine relationships. This approach is prone to errors because it relies heavily on language priors rather than systematic visual analysis. Human relational reasoning, by contrast, follows a more structured process: first locating and identifying relevant objects, then examining their interactions, and finally synthesizing visual evidence to draw conclusions about relationships. This multi-step approach allows for more careful consideration of visual information and reduces the likelihood of inference errors.
Inspired by this human reasoning process and drawing insights from Interleaved Modal Chain-of-Thought (ICoT) Gao et al. (2025), we propose ChainMPQ (Multi-Perspective Questions guided Interleaved Chain of Image and Text), a training-free framework that decomposes relational inference into manageable steps while maintaining coherent reasoning through accumulated multimodal memory. Our approach addresses the core challenge of relation hallucination by combining enhanced visual attention with systematic question decomposition and progressive multimodal reasoning.
Specifically, ChainMPQ first extracts subject and object keywords from the question and uses them to drive cross-attention. Next, relations are decomposed into three components, and multi-perspective questions are constructed by masking one component at a time. Finally, the interleaved chain is built by sequentially issuing the constructed questions, using previous answers as textual context and previous attention maps on all visual tokens to adjust later attention. This process enables the model to reason over prior textual and visual memories and progressively resolve the relation. We demonstrate the effectiveness of ChainMPQ on LLaVA-1.5 Liu et al. (2024b) and InstructBLIP Dai et al. (2023), where it consistently reduces relation hallucinations on relation-focused benchmarks. Our contributions are summarized as follows:
@@
    We introduce a subject–object–relation decomposition and generate multi-perspective questions that are closely centered on the three core elements of a relationship, encouraging the model to focus on each component.
@@ 2.1 Large Vision-Language Models (LVLMs)
While these models have achieved remarkable progress in visual-language understanding Lai et al. (2024); Laurençon et al. (2024); Li & Li (2025); Peng et al. (2024), they continue to face challenges of hallucinations, particularly in the recognition of relationships Zheng et al. (2024); Nie et al. (2024) between objects. Our work extends these models by specifically addressing the problem of relation hallucinations, employing a step-by-step reasoning process guided by both textual and visual information.
@@ 2.2 Relation Hallucination in LVLMs
Recent research has proposed several evaluation benchmarks aimed at assessing the handling of inter-object relations. For example, MMRel Nie et al. (2024) features over 10k question-answer pairs across multiple domains, specifically designed to evaluate spatial, action, and comparative relations. Additionally, R-Bench Wu et al. (2024) includes a diverse set of questions that balance perception and cognition tasks, exposing limitations in the relation reasoning capabilities of contemporary models. These benchmarks are valuable for systematically evaluating and enhancing the management of relation hallucinations. In this paper, we evaluate our methods using both MMRel and R-Bench.
@@ 3.5 Interleaved Text-Image Reasoning Chain
Extract keywords $K$ from $Q$; encode $V\leftarrow\mathrm{VisualEncoder}(I)$ and $X\leftarrow\mathrm{TextEncoder}(K)$;
@@
Q1 and Q2: localize $[S]$ and $[O]$ using $V^{\prime}$, update $\mathcal{T}$.
Q3 to Q5: ask relation-focused questions; build top-$k$ mask $M_{i}$ with $k\leq k_{\max}$,
apply bias scaled by $\lambda$, update $\mathcal{T}$ and $\mathcal{V}$.
@@
$\text{Attn}_{i}=\frac{1}{|T|\cdot K}\sum_{t\in T}\sum_{\ell=L-n}^{L-1}\text{Attn}^{(\ell)}[t,:]$ (2)
where T contains indices of keyword tokens, L is the total number of decoder layers, and $A^{(\ell)}$ represents the attention matrix at layer $\ell$.
@@
$\alpha=\lambda\cdot\mathit{Conf}_{\text{prev}_{i}}\qquad\text{Attn}_{i+1}=\operatorname{softmax}\!\left(\frac{QK^{\top}}{\sqrt{d_{k}}}+\alpha\cdot M_{i}\right)V$ (5)
where $\alpha$ is a confidence-based weight that increases with answer certainty.
@@ 4.1 Experimental Setup
Evaluated LVLMs. We evaluate the effectiveness of our method on two state-of-the-art open-source LVLMs: LLaVA-1.5-7B Liu et al. (2024b) and InstructBLIP-7B Dai et al. (2023). LLaVA-1.5 combines a CLIP visual encoder with Vicuna-7B for multimodal reasoning, while InstructBLIP-7B follows the BLIP-2 design, linking a frozen vision encoder to a language model through a Q-Former. These distinct architectures allow us to assess the generalization of our method across different design paradigms.
Dataset and Benchmarks. Commonly used benchmarks for visual question answering, such as POPE Li et al. (2023c) and CHAIR Rohrbach et al. (2018), primarily focus on object hallucinations. Since our task aims to alleviate relation hallucinations, we evaluate on two widely used benchmarks specifically designed for this purpose: MMRel Nie et al. (2024) and R-Bench Wu et al. (2024). The datasets details are illustrated in Appendix.
Baselines. In addition to a standard multimodal large model, we compare our approach with several methods specifically designed to mitigate relation hallucinations. These include Constraint-Aware Prompting Wu et al. (2025a), which relies on text-based prompting, and Detect-then-Calibrate Zheng et al. (2024), which calibrates the final output layer using hidden states from intermediate layers. The performance of these baseline methods is obtained by re-implementing their experiments on our chosen datasets and experimental setup using publicly available code.
4.2 Main Results
Overall Performance. Table 1 shows ChainMPQ consistently outperforms baselines across both models and benchmarks. On MMRel, ChainMPQ achieves 65.20% accuracy with LLaVA-1.5, improving over the best baseline by 1.7%. Similar improvements are observed with InstructBLIP (65.14% vs 64.52%). The method shows particularly strong precision gains (4.17% higher than the best baseline using LLaVA in R-Bench), indicating reduced false positive relation predictions while simultaneous gains in F1 further confirm that the method improves overall reliability without sacrificing recall.
Cross-Model Generalization. ChainMPQ demonstrates consistent improvements across different architectures (LLaVA vs InstructBLIP), suggesting the approach is model-agnostic rather than exploiting specific architectural features.
@@ 4.3 Ablation Study
4.3.1 Ablation study of core components.
We examine the contributions of ChainMPQ’s core components. To assess the role of "Text-guided Attention Enhancement", we remove the enhancement of visual tokens in subject and object regions, such that in the multimodal chain we use $V$ rather than $V^{\prime}$. To test the importance of the "Construction of Multi-perspective Aware Text Prompts", we replace the five constructed questions with a single relationship question(the fifth constructed question) as the intermediate guide. To evaluate the contribution of the "Interleaved Text-image Reasoning Chain", we remove the multimodal chain and only use the text context from the previous answer.
As shown in Table 3(a), the full ChainMPQ achieves an accuracy of 65.20%. Removing text-guided attention reduces accuracy by 1.14%, suggesting that enhancing subject- and object-related tokens provides limited but measurable benefit in reducing hallucinations. Omitting multi-perspective questions decreases accuracy by 3.68%, confirming their importance in guiding step-by-step inference, which forms the cornerstone of the chain. Eliminating the interleaved chain lowers accuracy by 3.08%, indicating that transferring visual memory plays an equally important role in the reasoning process. Despite these degradations, all three ablated variants still outperform the baseline LLaVA (59.02% accuracy), while the complete ChainMPQ demonstrates the strongest synergistic effect.
@@ 4.3.2 Sensitivity analysis of hyperparameters
We also analyze the sensitivity of $k_{\text{max}}$ , the maximum value for top-K in Eq. 3. We set $k_{\text{max}}$ to 5%, 10%, 30%, and 50% of the total number of patches, corresponding to approximately 10, 20, 70, and 120 tokens. We also vary the maximum bias weight parameter $\lambda$, which is the coefficient of $\alpha$ in Eq. 5, and test values of 3, 5, and 7. As shown in Figure 3(b), performance peaks at $\lambda=5$ and $k_{\text{max}}=20$, where the accuracy reaches 65.20%. When $k_{\text{max}}$ is too large, it using almost all image tokens, which dilutes focus on relevant regions and disrupts the attention mechanism. Conversely, a very small $k_{\text{max}}$ may omit critical information, especially for questions involving dispersed features. Similarly, a large $\lambda$ makes the model overly reliant on historical memory, disrupting normal attention propagation and reducing accuracy, while a small $\lambda$ yields insufficient bias and performs close to the original LLaVA baseline.
@@ 4.4 Case Study
(b) Case study spatial category
@@
To intuitively illustrate the ChainMPQ process, we present two real examples from MMRel, namely an action case and a spatial case, which together account for over 90% of the MMRel dataset Nie et al. (2024).
Figure 4 illustrates the complete processing pipeline of ChainMPQ. In the action case "Does a man stand on a surfboard in the image?", the baseline answers "yes", confusing the true relation "riding" with "standing". ChainMPQ first localizes the subject and the object with two guiding questions, surfaces "riding", and propagates this cue through textual context and attention transfer to produce the correct answer "no". In the spatial case "Is a chair to the left of a trash bin in the image?", the baseline is again incorrect. ChainMPQ retrieves the locations of the chair and the bin under the desk, constrains subsequent reasoning to the correct regions, and then verifies the directional relation to conclude that the chair is to the "right" of the bin. Across both cases, decomposition supports reliable localization, suppresses spurious language priors, and provides a disciplined path for relation verification that yields answers aligned with the image.
Figure 5 presents that across both action and spatial cases, ChainMPQ yields attention maps that are more sharply concentrated on task relevant regions, especially the subject and the object. This pattern indicates that the focused attention elicited by earlier related subquestions is propagated through the chain via textual context and attention transfer, so that when answering the original question the model attends more precisely to the pertinent areas. In both examples, ChainMPQ can suppress attention to irrelevant background areas and focus attention on the interaction between the subject and the target object, aligning the attention trajectory with the queried relation. The resulting focus correlates with corrected predictions and fewer contradictions between the image and the answer, which suggests improved grounding and more faithful relational inference.

### v2

@@ Abstract
While Large Vision-Language Models (LVLMs) achieve strong performance in multimodal tasks, hallucinations continue to affect their reliability. Among the three categories of hallucinations, which include object, attribute, and relation, relation hallucinations account for the largest proportion but have received the least attention. To address this challenge, we propose ChainMPQ (Multi-Perspective Questions guided Interleaved Text-image Reasoning Chain), a training-free method that improves relational inference in LVLMs by utilizing accumulated textual and visual memories. ChainMPQ first extracts subject and object keywords from the question to enhance the corresponding image regions. It then constructs multi-perspective questions that focus on the three core components of a relationship: the subject, the object, and the relation that links them. These questions are sequentially input to the model, with textual and visual memories from earlier steps providing supporting context for subsequent ones, thereby forming an interleaved chain of image and text that guides progressive relational reasoning. Experiments on multiple LVLMs and benchmarks show that ChainMPQ substantially reduces relation hallucinations, while ablation studies further validate the effectiveness of its three core modules.
@@ 1 Introduction
Although prior work has substantially reduced object and attribute hallucinations through preference optimization Wu et al. (2025a); Yang et al. (2025), contrastive decoding Wu et al. (2025c); Zhang et al. (2025), and intermediate-layer modification Jiang et al. (2024; 2025); Wang et al. (2024), relation hallucination, which accounts for nearly 40% of all hallucinations Zheng et al. (2024), remains a salient challenge. Existing efforts to mitigate relation hallucinations have made some progress: dataset-driven approaches Xie et al. (2025) focus on constructing high-quality training and fine-tuning data, prompt-engineering methods Wu et al. (2025a) employ constrained perception prompts, Detect-then-Calibrate Zheng et al. (2024) selectively adjust the final logits with intermediate layer, and Triplet Description Wu et al. (2025b) converts the image into triplets, allowing the model to answer questions based on these structured descriptions instead of the raw image.
Despite their promising results, existing methods targted at relation hallucinations treat relational reasoning as single-step inference, expecting models to simultaneously identify entities and determine relationships. This approach is likely to produce errors because it relies heavily on language priors rather than systematic visual analysis. Human relational reasoning, by contrast, follows a more structured process: first locating and identifying relevant objects, then examining their interactions, and finally synthesizing visual evidence to draw conclusions about relationships. This multi-step approach allows for more careful consideration of visual information and reduces the likelihood of inference errors.
Inspired by human reasoning process and drawing insights from Interleaved Modal Chain-of-Thought (ICoT) Gao et al. (2025), we propose ChainMPQ (Multi-Perspective Questions guided Interleaved Text-image Reasoning Chain), a training-free framework that decomposes relational inference into manageable steps while maintaining relevant reasoning through accumulated multimodal memory. Our approach addresses the core challenge of relation hallucination by combining systematic question decomposition with progressive multimodal reasoning.
Specifically, ChainMPQ first extracts subject and object keywords from the question and uses them to drive cross-attention. Next, relations are decomposed into three components, and multi-perspective questions are constructed by masking one component at a time. Finally, the interleaved chain is built by sequentially feeding the constructed questions into the model, using previous answers as textual context and earlier relevant visual tokens to adjust later attention maps. This process enables the model to reason over prior textual and visual memories and progressively resolve the relation. We demonstrate the effectiveness of ChainMPQ on four advanced and widely used models, where it consistently reduces relation hallucinations on relation-focused benchmarks. Our contributions are summarized as follows:
@@
    We introduce a subject–object–relation decomposition to generate multi-perspective questions, encouraging the model to focus on each core element of a relationship.
@@ 2.1 Large Vision-Language Models (LVLMs)
While these models have achieved remarkable progress in visual-language understanding Lai et al. (2024); laurençon2024buildingbetterunderstandingvisionlanguage; Li and Li (2025); Peng et al. (2024), they continue to face challenges of hallucinations, particularly in the recognition of relationships Zheng et al. (2024); Nie et al. (2024) between objects. Our work extends these models by specifically addressing the problem of relation hallucinations, employing a step-by-step reasoning process guided by both textual and visual information.
@@ 2.2 Relation Hallucination in LVLMs
Recent research has proposed several evaluation benchmarks aimed at assessing the handling of inter-object relations. For example, MMRel Nie et al. (2024) features over 10k question-answer pairs across multiple domains, specifically designed to evaluate spatial, action, and comparative relations. Additionally, R-Bench Wu et al. (2024) includes a diverse set of questions that balance perception and cognition tasks, exposing limitations in the relation reasoning capabilities of contemporary models. Tri-HE Wu et al. (2025b) further extends evaluation by using triplets to represent knowledge, which allows it to assess both object and relation hallucinations at the same time. These benchmarks are valuable for systematically evaluating and enhancing the management of relation hallucinations. For considerations of dataset size and usage frequency, we evaluate our method using MMRel and R-Bench, which are larger and more widely used.
@@ 3.5 Interleaved Text-Image Reasoning Chain
Extract keywords $K$ from $Q$;
Encode $V\leftarrow\mathrm{VisualEncoder}(I)$ and $X\leftarrow\mathrm{TextEncoder}(K)$;
@@
Q1 and Q2: localize $[S]$ and $[O]$ using $V^{\prime}$, update $\mathcal{T}$;
Q3 to Q5: ask relation-focused questions and build top-$k$ mask $M_{i}$ with $k\leq k_{\max}$;
Apply bias scaled by $\lambda$, update $\mathcal{T}$ and $\mathcal{V}$;
@@
$\text{Attn}_{i}=\frac{1}{|T|\cdot n}\sum_{t\in T}\sum_{\ell=L-n}^{L-1}\text{Attn}^{(\ell)}[t,:]$ (2)
where T contains indices of keyword tokens, L is the total number of decoder layers, and $\text{Attn}^{(\ell)}$ represents the attention matrix at layer $\ell$.
@@
$\alpha_{i}=\lambda\cdot\mathit{Conf}_{\text{prev}_{i}}\qquad\text{Attn}_{i+1}=\operatorname{softmax}\!\left(\frac{QK^{\top}}{\sqrt{d_{k}}}+\alpha_{i}\cdot M_{i}\right)V$ (5)
where $\alpha_{i}$ is a confidence-based weight that increases with answer certainty.
@@ 4.1 Experimental Setup
Evaluated LVLMs. We evaluate our method on four open-source LVLMs: LLaVA-1.5-7B Liu et al. (2024b), InstructBLIP-7B Dai et al. (2023), Qwen2.5-VL-7B Bai et al. (2025), and InternVL3.5-8B Wang et al. (2025). The first two models are widely used and represent classic LVLM designs. The last two models are more recent and show stronger performance in many multimodal tasks. LLaVA-1.5 combines a CLIP visual encoder with Vicuna-7B for multimodal reasoning, while InstructBLIP-7B follows the BLIP-2 design, linking a frozen vision encoder to a language model through a Q-Former. Qwen2.5-VL is built on the Qwen2.5 language model and uses a redesigned ViT architecture to provide strong visual understanding. InternVL3.5 integrates an InternViT visual encoder with a Qwen3 series language model to support robust image reasoning. These distinct architectures allow us to assess the generalization of our method across different design paradigms.
Dataset and Benchmarks. Commonly used benchmarks for visual question answering, such as POPE Li et al. (2023c) and CHAIR Rohrbach et al. (2018), primarily focus on object hallucinations. Since our task aims to alleviate relation hallucinations, we evaluate on two widely used benchmarks specifically designed for this purpose: MMRel Nie et al. (2024) and R-Bench (image-level) Wu et al. (2024). The datasets details are illustrated in Appendix.
@@
Qwen2.5-VL Vanilla 66.10 63.07 72.14 79.85 75.19 80.88
Prompting Wu et al. (2025a) 70.01 68.59 75.23 81.02 78.38 82.46
Calibrate Zheng et al. (2024) 71.36 67.20 76.28 80.98 77.05 81.52
CoT  Wei et al. (2023) 67.35 65.49 72.08 80.24 76.94 81.64
Ours 73.52 69.65 77.45 83.92 80.23 84.63
InternVL3.5 Vanilla 71.44 67.07 75.86 82.33 78.87 83.21
Prompting Wu et al. (2025a) 73.01 67.65 77.12 83.97 80.31 84.87
Calibrate Zheng et al. (2024) 74.45 71.46 78.64 82.65 79.82 83.75
CoT  Wei et al. (2023) 72.10 68.74 76.35 82.87 79.41 84.02
Ours 75.21 72.54 78.65 85.05 82.85 85.91
Baselines. In addition to a standard multimodal large model, we compare our approach with standard Chain-of-Thought prompting Wei et al. (2023) and several training-free methods specifically designed to mitigate relation hallucinations. These include Constraint-Aware Prompting Wu et al. (2025a), which relies on text-based prompting, and Detect-then-Calibrate Zheng et al. (2024), which calibrates the final output layer using hidden states from intermediate layers. We chose them because they represent two main ideas in this area: text-prompt-based reasoning and output calibration. Comparing with them allows us to show the benefit of our full text-image reasoning chain in a clear way. The performance of these baseline methods is obtained by re-implementing their experiments on our chosen datasets and experimental setup using the latest public code and their official instructions.
4.2 Main Results
Overall Performance. Table 1 shows ChainMPQ consistently outperforms baselines across both models and benchmarks. On MMRel, ChainMPQ achieves 65.20% accuracy with LLaVA-1.5, improving over the best baseline by 1.7%. Similar improvements are observed with InstructBLIP (65.14% vs 64.52%). On Qwen2.5-VL and InternVL3.5, which already have higher base performance, ChainMPQ also improves accuracy from 71.36% to 73.52% and from 74.45% to 75.21% respectively. The method shows particularly strong precision gains (4.17% higher than the best baseline using LLaVA in R-Bench), indicating reduced false positive relation predictions while simultaneous gains in F1 further confirm that the method improves overall reliability without sacrificing recall.
Cross-Model Generalization. ChainMPQ demonstrates consistent improvements across different architectures (LLaVA, InstructBLIP, Qwen-VL and InternVL), suggesting the approach is model-agnostic rather than exploiting specific architectural features.
4.3 Performance Experiment
To further improve the efficiency of ChainMPQ, we explored two optimization methods to achieve a trade-off between accuracy and latency.
  1. 1.
    Light1: We keep only Q1, Q2, and Q5. Since Q1 and Q2 don’t need to pass text or visual information to each other, they can run in parallel. When answering Q5, we use the answers of Q1 and Q2 as the text context, and their bias masks M1 and M2, constructed from the attention map, as the visual context.
  2. 2.
    Light2: We keep only Q3, Q4, and Q5, and all other settings remain the same. This means the reasoning chain starts from Q3 with no text or visual context, and the remaining steps (from Q3 to Q5) follow the same procedure described in Section 3.5.
The results in Table 2 show that the Light1 has the highest $\Delta\text{Acc}/\Delta\text{Time}$ , having a higher accuracy improvement in a shorter time, which means it achieves the best overall trade-off. Therefore, when accuracy is the top priority, we use the full ChainMPQ. When a balance between accuracy and latency is needed, Light1 is the preferred choice.
Table 2: Accuracy and latency comparison of different methods on LLaVA-1.5-7B.
Benchmark Method Acc (%) Time (s/sample) $\DeltaAcc / Δ\DeltaTime$
MMRel Vanilla 59.02 0.9
Full ChainMPQ 65.20 3.3 2.58
Light1 63.78 1.5 7.93
Light2 64.25 2.1 4.36
R-Bench Vanilla 71.23 0.9
Full ChainMPQ 76.04 3.3 2.00
Light1 74.50 1.5 5.45
Light2 75.08 2.1 3.21
@@ 4.4 Ablation Study
4.4.1 Ablation study of core components.
We examine the contributions of ChainMPQ’s core components. To assess the role of "Text-guided Attention Enhancement", we remove the enhancement of visual tokens in subject and object regions, such that in the multimodal chain we use $V$ rather than $V^{\prime}$.To test the importance of the "Construction of Multi-perspective Aware Text Prompts", we replace the five constructed questions with only the relationship question (Q5), meaning that the model only answers Q5 and then the original question. And The model answers the original question using the text context and visual bias produced when answering Q5. To evaluate the contribution of the "Interleaved Text-image Reasoning Chain", we remove the multimodal chain and only use the text context from the previous answer, which means that we keep the entire multi-turn text context, not only the last answer.
As shown in Table 3(a), the full ChainMPQ achieves an accuracy of 65.20%. Removing text-guided attention reduces accuracy by 1.14%, suggesting that enhancing subject- and object-related tokens provides limited but measurable benefit in reducing hallucinations. Omitting multi-perspective questions decreases accuracy by 3.68%, confirming their importance in guiding step-by-step inference, which forms the cornerstone of the chain. Eliminating the interleaved chain lowers accuracy by 3.08%, indicating that transferring visual memory plays an equally important role in the reasoning process. Despite these degradations, all three ablated variants still outperform the baseline LLaVA (59.02% accuracy), while the complete ChainMPQ demonstrates the strongest effect.
@@ 4.4.2 Sensitivity analysis of hyperparameters
We also analyze the sensitivity of $k_{\text{max}}$ , the maximum value for top-K in Eq. 3. We set $k_{\text{max}}$ to 5%, 10%, 30%, and 50% of the total number of patches, corresponding to approximately 10, 20, 70, and 120 tokens. We also vary the maximum bias weight parameter $\lambda$, which is the coefficient of $\alpha$ in Eq. 5, and test values of 3, 5, and 7. As shown in Figure 3(b), performance peaks at $\lambda=5$ and $k_{\text{max}}=20$, where the accuracy reaches 65.20%.When $k_{\text{max}}$ is too large, the model uses almost all image tokens, which weakens its focus on important regions and breaks the normal attention process. In contrast, when $k_{\text{max}}$ is too small, the model may miss key information, especially for questions that rely on scattered visual features. Similarly, if $\lambda$ is too large, the model becomes overly dependent on past memory, which disturbs attention propagation and reduces accuracy. If $\lambda$ is too small, the bias is weak and the performance is close to the original LLaVA baseline.
@@ 4.5 Case Study
To intuitively illustrate the ChainMPQ process, we present two real examples from MMRel, namely an action case and a spatial case, which together account for over 90% of the MMRel dataset Nie et al. (2024). Due to page limits, another case study of competitive type is provided in Appendix A.6.2.
@@
(b) Case study for spatial category
@@
Figure 4 illustrates the complete processing pipeline of ChainMPQ. In the action case "Does a man stand on a surfboard in the image?", the baseline answers "yes", confusing the true relation "riding" with "standing". ChainMPQ first localizes the subject and the object with two guiding questions, identifies the action “riding,” and propagates this cue through textual context and attention transfer to produce the correct answer "no". In the spatial case "Is there a chair left to a trash bin in the image?", the baseline is again incorrect. ChainMPQ retrieves the locations of the chair and the bin under the desk, constrains subsequent reasoning to the correct regions, and then verifies the directional relation to conclude that the chair is to the "right" of the bin. Across both cases, decomposition helps the model locate the correct regions, reduce false language priors, and follow a clear path to verify relations, leading to answers that aligned with the image.
Figure 5 presents that across both action and spatial cases, ChainMPQ produces attention maps that are more sharply concentrated on task-relevant regions, especially the subject and the object. This pattern indicates that the focused attention triggered by earlier related subquestions is propagated through the chain via textual context and attention transfer, so that when answering the original question the model attends more precisely to the related areas. In both examples, ChainMPQ can suppress attention to irrelevant background areas and focus attention on the interaction between the subject and the target object, aligning the attention trajectory with the queried relation. The resulting focus correlates with corrected predictions and fewer contradictions between the image and the answer, which suggests improved grounding and more faithful relational inference.